The mean is 60 with an SD of 15. What is the probability of selecting at random a score that falls between 55 and 65?
Answer #1

solution:-
given that mean = 60
standard deviation = 15

=>p(55 < x < 65) = p((55-60)/15 < z < (65-60)/15)

= p(-0.33 < z < 0.33

= p(z < 0.33) − p(z < −0.33)
= 0.6293 - 0.3707
= 0.2586